OpenStudy (anonymous):

Add. Simplify by collecting like radical terms if possible, assuming that all expressions under radicals represent non-negative numbers. √7a+4√63a^3

OpenStudy (saifoo.khan):

\[\Large \sqrt{7a}+4\sqrt{63a^3}\]

OpenStudy (saifoo.khan):

like that^

OpenStudy (anonymous):

correct

OpenStudy (saifoo.khan):

solving.

OpenStudy (saifoo.khan):

gotya!

OpenStudy (saifoo.khan):

\[\Large \sqrt{7a}(12a+1)\]

OpenStudy (anonymous):

that is the answer?

OpenStudy (saifoo.khan):

yesh! :)